# Navigating the Future of Federated Recommendation Systems with Foundation Models

### v1

@@ Abstract
In recent years, the integration of federated learning (FL) and recommendation systems (RSs), known as Federated Recommendation Systems (FRSs), has attracted attention for preserving user privacy by keeping private data on client devices. However, FRS faces inherent limitations such as data heterogeneity and scarcity, due to the privacy requirements of FL and the typical data sparsity issues of RSs. Foundation models (FMs), such as diffusion models (DMs) and large language models (LLMs), which are widely recognized for producing high-quality content in the image and NLP domains, focus on understanding and mimicking the underlying distribution of the training data. Unlike discriminative models, which learn the boundaries between categories, FMs aim to learn the entire probability distribution of the input data. Thus, the achievements of FMs inspire the design of FRS and suggest a promising research direction: integrating foundation models to address the above limitations. In this study, we conduct a comprehensive review of FRSs with FMs. Specifically, we: 1) summarise the common approaches of current FRSs and FMs; 2) review the challenges posed by FRSs and FMs; 3) discuss potential future research directions; and 4) introduce some common benchmarks and evaluation metrics in the FRS field. We hope that this survey provides the necessary background and guidance to explore this interesting and emerging topic.
@@ I Introduction
In the digital age, the exponential growth of information has created a need for systems that navigate, filter, and personalize data for individual users. Recommendation Systems (RS) nowadays become crucial tools for filtering online information and helping users discover products, content, and services that align with their preferences [1]. However, the systems traditionally rely heavily on centralized data collection and processing, posing significant privacy risks and operational bottlenecks. The importance of user privacy has never been greater, particularly with stringent data protection regulations such as the General Data Protection Regulation [2] in Europe, which emphasises the need to store user data on theirs local devices, instead of uploading it to central servers. As a novel approach to address these privacy concerns, Google introduced Federated Learning (FL) [3] as a framework designed to train models across decentralized devices, while keeping data localized. This paradigm shift in data processing leverages the computational capabilities of individual devices for local data analysis. Specifically, FL alternates between local model training at the user end and global parameter aggregation from these models on a central server. The integration of FL with RS becomes essential for safeguarding user privacy in recommendation services, which has given rise to the burgeoning field of Federated Recommendation Systems (FRS) [4]. In this field, typically each client consists of a single user’s device. Therefore, unless specified otherwise in this paper, the terms ’user,’ ’client,’ and ’device’ all refer to an individual user. FRS recently have shown promising results in many areas, such as service providing [5, 6], daily scheduling [7], driving planing [8, 9] and more, significantly impacting different facets of daily life. Similar to FL, FRS is required to keep user data local to protect user privacy. However, in most cases, each client typically contains only the data of one user’s accessed item, which is extremely small compared to the total number of items, creating a serious data sparsity problem. In addition, different users have different behaviours and preferences, which can lead to data heterogeneity. The presence of both issues can lead to sub-optimal models and reduced effectiveness.
@@
Figure 1: The foundation model can integrate information contained in data from various modalities during pre-training. The model can then be adapted for a variety of downstream tasks through adaptations such as prompting or fine-tuning.
@@
More recently, the emergence of a novel paradigm for the construction of artificial intelligence (AI) systems has garnered considerable interest in the wake of the remarkable success of ChatGPT [10] and Stable Diffusion [11] in the tasks of language understanding [12, 13, 14, 15] and image generation [16, 17], which we refer to as Foundation Models (FM) [18]. As shown in Fig. 1, this paradigm is built by using self-supervised optimization of training goals to determine how to update model parameters based on model predictions on the large amount of unlabelled training data. This process is referred to as pre-training. Language models, e.g. BERT [19]and RoBERT [20], are usually trained using the next token prediction goal, which refers to the extent to which the model is able to predict the next token in the sequence. One of the most successful examples of language models is ChatGPT, which is based on GPT-3.5. By training on large amounts of text data, it aligns the capabilities of large language models with human intent [21]. Vision models like ViT [22] are typically trained using either contrast learning or diffusion training targets. For contrast learning [23, 24], images are randomly augmented before evaluating the similarity of the model representations. For diffusion models [25, 26], noise is added to the images and the model is gradually de-noised by the target. There are also multi-modal training targets, some of which separate images and text during training, while others consider both. The foundation model, once trained, is used as a plug-in in combination with adaptations to achieve results for a wide variety of downstream tasks. CLIP [27] and DALL·E [28] are two multimodal AI models developed by OpenAI. CLIP focuses on understanding images through natural language, while DALL·E focuses on generating images based on text descriptions. While both models are trained on large datasets of images and text, CLIP is primarily used for image retrieval and classification, while DALL·E excels at generating new images that match text descriptions [29].
@@
Paper Organization. The reminder is structured to provide a thorough examination of the current research landscape, technical challenges, application cases, and future directions of FRS. The organization is as follows: Sec. II outlines the existing surveys within the field of FRS, and distinctly highlights our novel contributions, setting our survey apart from the current literature. Sec. III introduce foundational knowledge in the field of FRS and FM, including definitions, basic architectures, and current scenario classifications. Sec. IV introduces the potential applications and functions of FM according to the three classical phases of FRS. Sec. V delves into the main technical challenges faced by FRS with FM, including data heterogeneity, model aggregation difficulties, and privacy concerns, as well as current methods and their limitations in addressing these challenges. Sec. VII showcases applications of FRS across various domains, such as e-commerce, lifestyle and healthcare, discussing the outcomes and challenges of these real-world implementations. Sec. VIII summarizes the datasets and evaluation methods commonly used in FRS with FM, explaining how to quantify the recommendation performances. Sec. VI explores future research directions in FRS with FM, including unresolved issues and emerging research areas, providing guidance and inspiration for researchers in the field. Sec. IX concludes the survey with a summary of key findings and perspectives, offering insights into the trends and potential future of FRS.
@@ II Related Work and Contribution
In the field of FRSs, various studies have converged on methodologies, privacy preservation, and challenges, albeit with distinct focal points. Yang et al. [30] discuss the practical implementation and evaluation of federated recommendation systems, with a focus on system architecture and algorithmic efficiency. Alamgir et al. [31] provides a comprehensive overview of federated recommendation systems, highlighting techniques, prevailing challenges, and future directions. Javeed et al. [32] focus on the security and privacy concerns in personalized recommendation systems and propose targeted solutions to these challenges. Finally, Sun et al. [33] conduct a survey that compares existing federated recommendation research, highlighting the strengths and limitations of various approaches. All studies emphasize the critical importance of privacy protection and the challenges posed by data heterogeneity and model aggregation. This body of work enhances our understanding of FRSs and provides a foundation for future research to address their inherent challenges. Although a number of surveys [34, 35, 36, 37, 38, 39] focusing on FL with FM already exist, our paper is the first survey work in our current knowledge that focuses on combining FRS with FM.
@@ III Background and Preliminary
III-A Federated Recommendation System
Figure 3: A typical framework for a FRS.
@@ III-A Federated Recommendation System
FRS facilitate collaborative learning by pooling analysis and learning capabilities across multiple clients (e.g. users or locations), without the need for direct exchange of raw user data. Fig. 3 illustrates that this type of system typically involves three processes: client model update, communication, and global aggregation. The client model update allows users to train their models locally on their devices using their own data and then upload the updated intermediate parameters to the server. The server performs global aggregation on the parameters sent by all participants, integrating the unique information from each of them. The aggregated parameters are then distributed to the next round of participants, initiating a new training round. The process of uploading and distributing is collectively referred to as communication. These approaches effectively preserves privacy and security.
@@
III-A1 Taxonomy of FRS
  • Vertical FRS is applied when the feature spaces differ but the sample spaces are similar. This scenario is common when different entities possess different types of data about the same users. An example would be a bank and an online retailer holding distinct perspectives on the same customers, i.e., the bank has credit history while the retailer has shopping history. Vertical FRS [45, 46, 47, 48] trains recommendation models by securely integrating different data features between parties, utilizing richer user information for more accurate personalized recommendations and ensuring data privacy and security.
  • Transfer FRS utilizes principles of transfer learning [49] to transfer knowledge from one domain to another. This type is suited for scenarios with significant differences in both feature and sample spaces across parties. Transfer learning allows the system to leverage data and knowledge from the source domain to enhance recommendation performance in the target domain, even when the target domain lacks sufficient data for independent model training. Transfer FRS [50, 51] is particularly beneficial for emerging markets or user groups with limited data, drawing insights from related domains.
@@ III-A1 Taxonomy of FRS
  • Centralized FRS operates within a FL framework where a central server orchestrates the learning process. User devices (or clients) locally compute updates based on their data and send these updates to the server. The server aggregates these updates to improve the global model, which is then distributed back to the users. While this approach improves privacy by not requiring the sharing of raw data, it still relies on a central authority to manage the model. It effectively addresses privacy concerns by allowing the model to learn from decentralized data sources without centralizing the data itself. This architecture is typically used to overcome data silo issues and improve the performance of RSs without compromising user privacy and data security. Due to its simplicity and intuitive nature, this architecture has become the dominant framework within the field of FRS. There is a great deal of work [52, 43, 44] based on it currently.
  • Semi-decentralized FRS introduces an intermediate layer between the central server and the users, such as edge servers or devices that can perform additional computational or storage tasks. This setup aims to reduce the communication overhead and latency associated with sending updates to a central server, especially in large-scale applications. A specific example is the Semi-decentralized Federated Ego Graph Learning (SemiDFEGL) framework [53], which improves scalability and reduces communication costs by introducing new device-to-device collaborations. It augments local subgraphs with predicted interacted item nodes to exploit high-order collaborative information between users and items in a privacy-preserving manner, which is particularly useful for recommendations based on collaborative filtering and graph neural networks.
  • Decentralized FRS, which employs peer-to-peer communication architecture [54], distributes the learning process completely to all participating devices, without the need for a central server for model aggregation. This approach maximizes privacy and data ownership, but presents challenges in coordinating model updates and ensuring model convergence. Zheng et al. [55] proposed a FRS model called DGREC by adopting a decentralized graph neural network, which constructs a local intra-item hypergraph and a global inter-user graph for each user, allowing users to freely choose whether to disclose their interactions. Li et al. [56] introduced DFedRec, which uses a privacy-aware client-level structured graph to share model parameters only with relevant neighboring users, thereby reducing communication costs and protecting user privacy.
@@ III-B Foundation Model
A foundation model is defined as any model trained on extensive data (typically using large-scale self-supervised learning) that is capable of adapting to a wide range of downstream tasks, for instance, through fine-tuning.
In recent years, the scale and scope of FM have greatly expanded our imagination of potential applications. Such models typically have billions or even trillions of parameters, allowing them to learn more complex patterns and knowledge. They can be adapted to new tasks through fine-tuning or zero-sample learning. For instance, the GPT-3 model [57], comprising 175 billion parameters, is capable of adapting and perform various tasks with the aid of natural language prompts, despite the fact that a significant proportion of these tasks have not been explicitly trained. FM is distinguished by two key characteristics: 1) Emergence, which refers to the implicit induction of system behavior from examples, as opposed to explicit design; and 2) Homogenization, which indicates that the method of constructing machine learning systems tends to be unified across a wide range of applications. Although FM is based on standard deep learning and transfer learning, their scale brings new emergent abilities, and their effectiveness in many tasks has motivated homogenization. While homogenization provides strong leverage, it also requires caution, as defects in foundation models can be inherited by all downstream adapted models. Moreover, despite the broad deployment prospects of foundation models, due to their emergent properties, our understanding of how they work, when they fail, and what they can actually do is still quite limited.
@@ III-B1 Taxonomy of FM
  • Language FM: These types of FM mainly deal with textual data and are trained on large textual datasets to understand and generate language. They are very effective in natural language processing (NLP) tasks such as machine translation, text summarization, and sentiment analysis. Typical language models include BERT [19], GPT-3 [57], and T5 [58], which typically use deep transformer architectures that are able to capture complex linguistic regularities and demonstrate excellent performance on multilingual tasks.
@@
  • Vision FM: Vision FM focuses on processing and understanding image data. These models are able to perform tasks such as object recognition, image segmentation, and visual reasoning by learning large amounts of image data. For example, DINOv2 [59] and SAM [60] are basic models trained specifically for visual tasks, and using self-supervised learning methods, these models learn valid visual representations without labelling the data.
@@
  • Multimodal FM: Multimodal models can simultaneously process and understand multiple types of data, such as text and images. Such models can excel in cross-modal tasks such as image captioning and visual quizzing by integrating information from different modalities. CLIP [27] and DALL·E [28] are representative of such models that are able to understand the relationship between an image and its corresponding textual descriptions, demonstrating flexibility and robustness in handling multiple data types.
Functionally, FMs are generally categorized into two types [18, 61]:
@@
  • Discriminative FM: The main task of discriminative FM is to distinguish or predict specific outputs from given input data. These models are generally based on the BERT family and are more concerned with learning decision boundaries from the data to perform classification, regression, or other predictive tasks.
@@
  • Generative FM: The core goal of generative FM is to learn the distribution of the data and be able to generate new data samples. These models, such as GPT-3, DALL·E, etc., are able to capture the underlying structure of the data and thus generate new instances that are similar but different from the training data. Generative FM has a wide range of applications in areas such as text generation and image generation.
III-B2 Adapter

### v2

@@ Abstract
In recent years, the integration of federated learning (FL) and recommendation systems (RS), known as Federated Recommendation Systems (FRS), has attracted attention for preserving user privacy by keeping private data on client devices. However, FRS faces inherent limitations such as data heterogeneity and scarcity, due to the privacy requirements of FL and the typical data sparsity issues of RSs. Models like ChatGPT are empowered by the concept of transfer learning and self-supervised learning, so they can be easily applied to the downstream tasks after fine-tuning or prompting. These models, so-called Foundation Models (FM), fouce on understanding the human’s intent and perform following their designed roles in the specific tasks, which are widely recognized for producing high-quality content in the image and language domains. Thus, the achievements of FMs inspire the design of FRS and suggest a promising research direction: integrating foundation models to address the above limitations. In this study, we conduct a comprehensive review of FRSs with FMs. Specifically, we: 1) summarise the common approaches of current FRSs and FMs; 2) review the challenges posed by FRSs and FMs; 3) discuss potential future research directions; and 4) introduce some common benchmarks and evaluation metrics in the FRS field. We hope that this position paper provides the necessary background and guidance to explore this interesting and emerging topic.
@@ I Introduction
In the digital age, the exponential growth of information has created a need for systems that navigate, filter, and personalize data for individual users. Recommendation Systems (RS) nowadays become crucial tools for filtering online information and helping users discover products, content, and services that align with their preferences [1]. However, the systems traditionally rely heavily on centralized data collection and processing, posing significant privacy risks and operational bottlenecks. The importance of user privacy has never been greater, particularly with stringent data protection regulations such as the General Data Protection Regulation (GDPR) [2] in Europe, which emphasises the need to store user data on theirs local devices, instead of uploading it to central servers. As a novel approach to address these privacy concerns, Google introduced Federated Learning (FL) [3] as a framework designed to train models across decentralized devices, while keeping data localized. This paradigm shift in data processing leverages the computational capabilities of individual devices for local data analysis. Specifically, FL alternates between local model training at the client and global parameter aggregation from these models on a central server. The integration of FL with RS becomes essential for safeguarding user privacy in recommendation services, which has given rise to the burgeoning field of Federated Recommendation Systems (FRS) [4]. In this field, typically each client consists of a single user’s device. Therefore, unless specified otherwise in this paper, the terms ’user,’ ’client,’ and ’device’ all refer to an individual user. FRS recently have shown promising results in many areas, such as service providing [5, 6], daily scheduling [7], driving planing [8, 9] and more, significantly impacting different facets of daily life. Similar to FL, FRS is required to keep user data local to protect user privacy. However, in most cases, each client typically contains only the data of one user’s accessed item, which is extremely small compared to the total number of item set, creating a serious data scarcity problem. In addition, different users have different behaviours and preferences, which can lead to data heterogeneity. The presence of both issues can lead to sub-optimal models and reduced effectiveness.
@@
Figure 1: The foundation model can integrate information contained in data from various modalities during pre-training. The model can then be adapted for a variety of downstream tasks through adapters such as prompting or fine-tuning.
@@
More recently, the emergence of a novel paradigm for the construction of artificial intelligence (AI) systems has garnered considerable interest in the wake of the remarkable success of ChatGPT [10] and diffusion models [11, 12, 13] in the tasks of language understanding [14, 15, 16, 17] and image generation [12, 13], which we refer to as Foundation Models (FM) [18]. As shown in Fig. 1, this paradigm is built by using self-supervised optimization of training goals to determine how to update model parameters based on model predictions on the large amount of unlabelled training data. This process is referred to as pre-training. Then, FM can utilize a technique, called Adaptation, to turn the general FM into a task-specific one. This process is termed as Fine-tuning or Prompting. Language models, e.g. BERT [19]and RoBERTa [20], are usually trained using the next token prediction goal, which refers to the extent to which the model is able to predict the next token in the sequence. One of the most successful examples of language models is ChatGPT, which is based on GPT-3.5. By training on large amounts of text data, it aligns the capabilities of large language models (LLM) with human intent [18]. Vision models like ViT [21] are typically trained using either contrast learning or diffusion training targets. For contrastive learning [22], images are randomly augmented before evaluating the similarity of the model representations. For diffusion models [11, 12, 13], noise is added to the images and the model is gradually de-noised by the target. There are also multi-modal training targets, some of which separate images and text during training, while others consider both. The foundation model, once trained, is used as a plug-in in combination with adaptations to achieve results for a wide variety of downstream tasks. CLIP [23] and DALL·E [24] are two multi-modal AI models developed by OpenAI. CLIP focuses on understanding images through natural language, while DALL·E focuses on generating images based on text descriptions. While both models are trained on large datasets of images and text, CLIP is primarily used for image retrieval and classification, while DALL·E excels at generating new images that match text descriptions [25].
@@ II Related Work and Contribution
In the field of FRSs, various studies have converged on methodologies, privacy preservation, and challenges, albeit with distinct focal points. Yang et al. [26] discuss the practical implementation and evaluation of FRS, with a focus on system architecture and algorithmic efficiency. Alamgir et al. [27] provides a comprehensive overview of FRS, highlighting techniques, prevailing challenges, and future directions. Javeed et al. [28] focus on the security and privacy concerns in personalized recommendation systems and propose targeted solutions to these challenges. Finally, Sun et al. [29] conduct a survey that compares existing federated recommendation research, highlighting the strengths and limitations of various approaches. All studies emphasize the critical importance of privacy protection and the challenges posed by data heterogeneity and model aggregation. This body of work enhances our understanding of FRS and provides a foundation for future research to address their inherent challenges. Although a number of study focusing on combining FM with FL [30, 31, 32, 33, 34, 35, 36] or RS [37, 38] already exist, our paper is the first work to explore the integration of FM into FRS.
@@ III Background and Preliminary
III-A Federated Recommendation System
Figure 2: A typical framework for FRS.
@@ III-A Federated Recommendation System
FRS facilitate collaborative learning by pooling analysis and learning capabilities across multiple clients (e.g. users or locations), without the need for direct exchange of raw user data. Fig. 2 illustrates that this type of system typically involves three processes: 1) client model update, 2) communication, and 3) global aggregation. The client model update allows users to train their models locally on their devices using their own data and then upload the updated intermediate parameters, e.g. the model gradients or parameters can be shared, to the server. The server then performs global aggregation on the parameters sent by all participants, integrating the unique information from each of them. The aggregated parameters are then distributed to the next round of participants, initiating a new training round. The process of uploading and distributing is collectively referred to as communication. These approaches effectively preserves privacy and security.
III-A1 Taxonomy of FRS
  • Vertical FRS is applied when the feature spaces differ, but the sample spaces are similar, as shown in Fig. 3b. This scenario is common when different parties possess different types of data about the same users. An example would be a bank and an online retailer holding distinct perspectives on the same customers, i.e., the bank has credit history while the retailer has shopping history. One challenge is that how to align the identities of customers across different parties. Vertical FRS [42, 43, 44, 45] trains recommendation models by securely integrating different data features between parties, utilizing richer user information for more accurate personalized recommendations and ensuring data privacy and security.
  • Transfer FRS utilizes principles of transfer learning [46] to transfer knowledge from one domain to another, as shown in Fig. 3c. This type is suited for scenarios with significant differences in both feature and sample spaces across parties. Transfer learning allows the system to leverage data and knowledge from the source domain to enhance recommendation performance in the target domain, even when the target domain lacks sufficient data for independent model training. Transfer FRS [47] is particularly beneficial for emerging markets or user groups with limited data, drawing insights from related domains.
@@ III-A1 Taxonomy of FRS
  • Centralized FRS operates within a FL framework where a central server orchestrates the learning process as shown in Fig. 5a. Clients locally compute updates based on their data and send these updates to the server. The server aggregates these updates to prompt essential information sharing, which is then distributed back to the users. It effectively addresses privacy concerns by allowing the model to learn from decentralized data sources without centralizing the data itself. While this approach improves privacy by not requiring the sharing of raw data, it still relies on a central authority to manage the model. This architecture is typically used to overcome data silo issues [48] and improve the performance of RS without compromising user privacy and data security. Due to its simplicity and intuitive nature, this architecture has become the dominant framework within the field of FRS. There is a great deal of work [49, 39, 40, 41, 47] based on it currently.
  • Semi-decentralized FRS shown in Fig. 5b introduces an intermediate layer between the central server and the users, such as edge servers or devices that can perform additional computational or storage tasks. This setup aims to reduce the communication overhead and latency associated with sending updates to a central server, especially in large-scale applications. A specific example is the SemiDFEGL framework [50], which improves scalability and reduces communication costs by introducing new device-to-device collaborations. It augments local subgraphs with predicted interacted item nodes to exploit high-order collaborative information between users and items in a privacy-preserving manner, which is particularly useful for recommendations based on collaborative filtering and graph neural networks.
  • Decentralized FRS, which employs peer-to-peer communication architecture [51], distributes the learning process completely to all participating devices, without the need for a central server for model aggregation, as shown in Fig. 5c. This approach maximizes privacy and data ownership, but presents challenges in coordinating model updates and ensuring model convergence. Zheng et al. [52] proposed a FRS model called DGREC by adopting a decentralized graph neural network, which constructs a local intra-item hypergraph and a global inter-user graph for each user, allowing users to freely choose whether to disclose their interactions. Li et al. [53] introduced DFedRec, which uses a privacy-aware client-level structured graph to share model parameters only with relevant neighboring users, thereby reducing communication costs and protecting user privacy.
@@ III-B Foundation Model
A foundation model is defined as any model trained on extensive data (typically using large-scale self-supervised learning) that is capable of adapting to a wide range of downstream tasks, for instance, through fine-tuning or prompting.
In recent years, the scale and scope of FM have greatly expanded our imagination of potential applications. Such models typically have billions or even trillions of parameters, allowing them to learn more complex patterns and knowledge. They can be adapted to new tasks through fine-tuning or prompting [54, 55]. For instance, the GPT-3 model [56], comprising 175 billion parameters, is capable of adapting and perform various tasks with the aid of natural language prompts, despite the fact that a significant proportion of these tasks have not been explicitly trained. FM is distinguished by two key characteristics [18]: 1) Emergence, which refers to the implicit induction of system behavior from examples, as opposed to explicit design; and 2) Homogenization, which indicates that the method of constructing machine learning systems tends to be unified across a wide range of applications. Although FM is based on self-supervised learning and transfer learning, their scale brings new emergent abilities, and their effectiveness in many tasks has motivated homogenization, e.g., many FMs have evolved based on transformer architectures [19, 20, 21, 56]. While homogenization provides strong leverage, it also requires caution, as defects in foundation models can be inherited by all downstream adapted models. Despite the broad deployment prospects of foundation models, due to their emergent properties, our understanding of how they work, when they fail, and what they can actually do is still quite limited [18], but there has been quite a bit of work [57, 58, 59] that attempts to explain this phenomenon. Moreover, OpenAI proposes a Scaling Law for LLM [60], suggesting that the performance of a model is governed by a power-law relationship involving computational power, model parameters, and data size. This law also applies to downstream tasks [61]. Therefore, understanding and leveraging this power-law relationship when using FM can aid in more efficiently allocating resources and optimizing model performance.
@@ III-B1 Taxonomy of FM
  • Language FM: These types of FM mainly deal with textual data and are trained on large textual datasets to understand and generate language [19, 20]. They are very effective in natural language processing (NLP) tasks such as machine translation [62], search [63, 64], and sentiment analysis [65]. Typical language models include BERT [19], GPT-3 [56], and LLama [66], which typically use deep transformer architectures that are able to capture complex linguistic regularities and demonstrate excellent performance on multilingual tasks.
@@
  • Vision FM: Vision FM focuses on processing and understanding image data. These models are able to perform tasks such as object recognition [21, 67, 68], image segmentation [69, 70, 71], and visual reasoning [72, 73] by learning large amounts of image data. For example, DINOv2 [74] and SAM [69] are basic models trained specifically for visual tasks, and using self-supervised learning methods, these models learn valid visual representations without labelling the data.
@@
  • Multi-modal FM: Multi-modal models can simultaneously process and understand multiple types of data, such as text and images. Such models can excel in cross-modal tasks such as image captioning and visual quizzing [75, 76, 77] by integrating information from different modalities. CLIP [23] and DALL·E [24] are representative of such models that are able to understand the relationship between an image and its corresponding textual descriptions, demonstrating flexibility and robustness in handling multiple data types.
Functionally, FMs are generally categorized into two types [18, 78, 38]:
@@
  • Discriminative FM: The main task of discriminative FM is to distinguish or predict specific outputs from given input data. These models are generally based on the BERT family and are more concerned with learning decision boundaries from the data to perform tasks like classification [79, 80] and regression [81].
@@
  • Generative FM: The core goal of generative FM is to learn the distribution of the data and be able to generate new data samples [25]. These models, such as GPT-3, DALL·E, etc., are able to capture the underlying structure of the data and thus generate new instances that are similar but different from the training data. Generative FM has a wide range of applications in areas such as the generation of text [82] and images [12, 13].
III-B2 Adaptation